# From underground natural gas to hydrogen storage in fractured reservoir rock : comparing relative permeabilities for hydrogen versus methane and nitrogen

### v1

@@ 3.2 Saturation and fluid distribution
To further explore the cause of the similarities and differences between the different relative permeability measurements, we investigated the average saturation and fluid distribution in the fracture space. The slice-averaged 1-D saturation profiles for hydrogen, methane, and nitrogen experiments at different fractional flows are shown in Figures 7(a), 7(b), and 8, respectively. As the fractional flow of gas (fg) increases, the slice-averaged brine saturation (Sw) across the sample decreases for all gases, indicating that more gas is occupying the fracture space. For hydrogen and methane, saturation varies between 0.3 and 1, while for nitrogen it ranges from 0.27 to 0.99, suggesting a slightly lower average saturation for nitrogen. It results from the fracture geometry being highly heterogeneous with variable aperture distribution and roughness as shown in Figure 2(b). The dry scan, Figure 1 reveals secondary fractures branching from the centre of the sample, contributing to the intricate fracture network. This complexity affects the distribution of fluid saturation across the sample, influencing gas displacement efficiency and the overall flow dynamics.
@@
At lowest fractional flows fg = 0.1, the saturation profiles illustrated in Figure 9(a) shows noticeable differences, with brine saturation for hydrogen being 5 higher than for methane and 10 higher than for nitrogen. This variation primarily results from hydrogen’s lower viscosity, which leads to less efficient invasion of the fractures, ultimately resulting in lower gas saturation and higher wetting-phase saturation. Although the fluid properties of hydrogen and methane at 10 MPa and 22°C are similar, nitrogen, in its supercritical state, demonstrates distinctly different flow behavior. The higher viscosity of nitrogen facilitates more efficient invasion of the fracture network, resulting in higher gas saturation. Nitrogen is observed to drain more efficiently compared to hydrogen and methane in porous media, as also reported by Higgs et al. [36]. Despite the interfacial tension (IFT) of nitrogen (0.07 N m-1) and hydrogen (0.069 N m-1) being comparable at 10 MPa, the differences in viscosity significantly influence wetting-phase saturation at onset of drainage. This underscores the finding that at low fractional flows, the viscosity ratio plays a more pronounced role in gas displacement behavior. The fluid distribution in the fracture at fg = 0.6, as shown in Figure 10, indicates that nitrogen has the highest gas fraction (Sw = 0.76) in the fracture, followed by methane (Sw = 0.8), and then hydrogen (Sw = 0.76).
@@
At higher gas fractional flows (fg = 0.8 and 1) the saturation converges for all gases as shown in Figure 9(b). At these higher gas fractional flows, the capillary number decreases resulting in more connected gas which diminishes the impact of the viscosity ratio. Consequently, similar saturations are observed at fg = 1, representing endpoint saturation. Although the nitrogen case has a significantly higher Ca (1.2 x 10-6) than hydrogen and methane (7.0 x 10-6 and 1.5 x 10-6 respectively), the end point saturations remain consistent for all gases at Sw = 0.6. This suggests that the effect of capillary number on overall saturation is relatively minor under the tested conditions.
@@
The fluid distribution for the hydrogen-brine system in the fracture exhibits signs of intermittency at higher water saturations (lower ), as illustrated in Figure 11 for . Subtle differences in gas distribution are observed between two scans taken 30 minutes apart after reaching steady-state pressure. Several locations are identified by coloured circles: Circle 1 (yellow) marks a region where gas disappears in the subsequent scan, while Circles 2 (green) indicate location where gas reappears. These fluctuations in the gas phase correspond to capillary events and are also reflected in the measured differential pressure. In contrast, the nitrogen-brine system exhibits fewer fluctuations in saturation, suggesting a more stable flow pattern. A more detailed analysis of pressure fluctuations, intermittency, and their impact on fluid distribution will be discussed in Section 3.3. Along the length of the fracture, slice-averaged saturation reaches Sw = 1 for many slices, indicating a disconnected gas phase. Figure 12 presents the 1-D slice-averaged saturation profiles for these two subsequent scans revealing significant fluctuations in saturation within this period, emphasizing dynamic capillary processes occurring in the fracture. These processes are not fully captured in terms of intermediate gray values due to the scanner’s limited temporal resolution. As the gas fractional flow increases to fg = 0.8, the gas phase becomes more connected throughout the fracture, which reduces intermittency and enhances overall gas saturation as demonstrated by the 3D volume rendering of fluid distribution in Figure 13. Intermittency significantly impacts relative permeability, resulting in notably low gas permeabilities, particularly for hydrogen and methane as discussed in Section 3.1.
@@
We note here that the observed intermittency may also be influenced by the dissolution of hydrogen and methane into the brine during the experiment, as the injected gas and brine were not pre-equilibrated, a phenomenon previously described by [33]. However, the reappearance of gas in the fracture downstream indicates that intermittency is a genuine effect and not solely attributed to dissolution. The solubilities of hydrogen, methane, and nitrogen in water are 0.8 mg L-1, 22.7 mg L-1, and 10 mg L-1, respectively. Hydrogen’s significantly lower solubility, yet comparable phase disconnectivity, suggests that dissolution effects are likely minimal.
@@ 3.3 Pressure analysis
The lower gas relative permeabilities for similar global water saturation suggest higher phase interference in the fracture. Saturation and fluid distribution in the fracture, based on scans taken 30 minutes apart, indicate the presence of intermittency, where Haines jumps and snap-off events occur. Since the scan duration is 16 minutes and these events appear to happen much more rapidly, we observed pressure data recorded at a much higher time resolution of 1 second. The fluctuations in differential pressure measurements from the transducer across the rock samples provide insights into the timescale and the fluctuations that correspond to snap-off and Haines jumps occurring during two-phase displacements [52, 51]. Differential pressure measurements at over a 30-minute period for H2, CH4, and N2 are presented in Figure 14. These data reveal fluctuations over varied timescales for each gas, underscoring distinct dynamics associated with each gas phase. There are notable differences in the amplitude and frequency of these oscillations, particularly between H2 and CH4 on one hand and N2 on the other. The pressure amplitude of fluctuations is greater for hydrogen and methane compared to nitrogen, with the fluctuations occurring over timescales of approximately 150 seconds for hydrogen, 30 seconds for methane, and one second for nitrogen. The pressure fluctuations within the pores result in intermittent pore filling, occurring over a broad range of timescales. At different fractional flows, the timescales of these fluctuations vary, reflecting changes in the connectivity of the gas phase. The additional pressure data and preliminary analysis, including the Gaussian distribution of fluctuations and red noise, are provided in Supplementary Information S2.
@@ 3.4 Sensitivity analysis
As discussed in previous sections, relative permeabilities at reservoir conditions in fractures are strongly influenced by intermittency, leading to significant fluctuations in pressure and saturation. To assess the sensitivity of the hydrogen-brine system under these conditions, we repeated the experiment under identical flow conditions. Overall, qualitative trends in the relative permeability-saturation curves remained consistent across both runs, as shown in Figure 15. The water permeabilities () were quantitatively similar in both runs. Gas permeability, while consistently low across runs, exhibited slight differences, likely due to minor pressure fluctuation variations. Since gas permeabilities are very low, even slight variations in pressure fluctuations can result in differences in the measured relative gas permeability. The pressure fluctuations for these two runs at are shown in Figure 15(b). The data clearly exhibits cyclic fluctuations, with the timescales of these fluctuations differing between the two runs: in the first run, the period was 0.5 minutes, while in the repeated run, the period extended to 2.5 minutes. These variations in fluctuation timescales may contribute to slight differences in the calculated relative gas permeabilities. Such differences could arise due to slight changes in fracture morphology, potentially occurring during pressurization, as the fracture is sensitive to applied stress. Over the course of running different experiments presented in this paper , the rock sample underwent pressurization-depressurization cycles, potentially impacting fracture morphology. This is reflected in the observed decrease in absolute permeability before the experimental runs, from 23 D to 8 D. Despite these slight variations, the overall behavior remains consistent across both experimental runs, confirming that our workflow is robust and validating the results presented in this study. The differential pressure data for other fractional flows are provided in Supplementary Information S3.
@@ 4 Discussion
Direct measurements of relative permeability for hydrogen compared to other gases in fractured rock samples are currently absent from the literature. Here, we compare our findings with existing experimental data and numerical models for different fluid systems in rough fractures, as shown in Figure 16. Compared to previous studies, the hydrogen gas relative permeability in our system is notably lower, indicating interference effects that are stronger than predicted by existing models such as [27], Gong et al. [29], and Bertels et al. [53]. The limited available experimental data were collected under different flow conditions and using different fluids, making direct comparisons challenging. The fluid pairs used in various studies from the literature are summarized in Supplementary Information S4. The nitrogen permeability aligns more closely with the model proposed by Gong et al. [29]; however, the impact of fracture geometry could not be fully resolved based on these comparisons. The nitrogen-brine system relative permeabilities were also compared to models from the literature, with detailed discussions provided in the Supplementary Information S4. In contrast, water permeability in our results aligns more closely with the model suggested by Piri and Karpyn [54]. Importantly, fracture morphology also influences these differences, so while such comparisons are informative, they are not definitive. Controlled experiments and models that account for specific fracture morphology in hydrogen-brine systems are essential for more conclusive insights.
@@
The lower hydrogen gas permeability observed in our fracture study is consistent with trends seen in porous media; thus, we also compared our measurements with direct hydrogen permeability data from porous media studies in Berea sandstone [32, 33] and Bentheimer sandstone [36], shown in Figure 17. Despite marked differences in sample geometry and flow conditions, the results remain qualitatively comparable, as all experiments were conducted at low capillary numbers (Cat). In porous media, the relative permeability curves from Higgs et al. [36] exhibit closer alignment for hydrogen, methane, and nitrogen. However, our fracture study reveals greater differences in both wetting- and non-wetting-phase permeabilities. This divergence can likely be attributed to the increased intermittency in the fracture, resulting from the higher pressure conditions (100 bar) in our study, where the interfacial tension (IFT) between hydrogen and methane differs more significantly, and nitrogen exists in a supercritical state. In contrast, the experiments by Higgs et al. [36], conducted at 20 bar, were performed under conditions where the fluid properties of these gases were more similar. Additionally, our study shows a much higher irreducible saturation, exemplifying the dependence of capillary-dominated relative permeability curves on flow rate (injection rate) and geometry (porous vs. fractured). When comparing our hydrogen permeability results to those from Boon and Hajibeygi [33], who also conducted high-pressure experiments (100 bar) at 18∘C under similar conditions, hydrogen and water permeabilities were generally higher across most saturations for porous media, though the general trend remains consistent. These comparisons highlight the increased complexity of fluid interactions in fractures compared to those observed in porous media.
@@ 5 Conclusions
In conclusion, this study provides critical insights into hydrogen behavior in fractured rock, supporting efforts to optimize underground hydrogen storage. By highlighting the limitations of using nitrogen as a proxy for hydrogen, our findings emphasize the need for hydrogen-specific research to develop effective storage strategies and safety protocols in subsurface reservoirs.
Future work should focus on measuring hydrogen-brine relative permeability in fracture samples with varied aperture sizes and roughness. Such studies will be valuable in understanding the impact of fracture geometry on flow behavior, ultimately enhancing the accuracy of upscaled fracture network models for reservoir-scale applications.

### v2

@@ 3.2 Saturation and fluid distribution
To further explore the cause of the similarities and differences between the different relative permeability measurements, we investigated the average saturation and fluid distribution in the fracture space. The slice-averaged 1-D saturation profiles for hydrogen, methane, and nitrogen experiments at different fractional flows are shown in Figures 5(a), 5(b), and 5(c), respectively. As the fractional flow of gas () increases, the slice-averaged brine saturation (Sw) across the sample decreases for all gases, indicating that more gas is occupying the fracture space. For hydrogen and methane, saturation varies between 0.3 and 1, while for nitrogen it ranges from 0.27 to 0.99, suggesting a slightly lower average saturation for nitrogen-brine system. It results from the fracture geometry being highly heterogeneous with variable aperture distribution and roughness as shown in Figure 1(d).
@@
At lowest fractional flows = 0.1, the saturation profiles illustrated in Figure 6(a) shows noticeable differences, with brine saturation for hydrogen being 5 higher than for methane and 10 higher than for nitrogen. This variation primarily results from hydrogen’s lower viscosity, which leads to less efficient invasion of the fractures, ultimately resulting in lower gas saturation and higher wetting-phase saturation. Although the fluid properties of hydrogen and methane at 10 MPa and 22∘C are similar, nitrogen, in its supercritical state, demonstrates distinctly different flow behavior. The higher viscosity of nitrogen facilitates more efficient invasion of the fracture network, resulting in higher gas saturation. Nitrogen is observed to drain more efficiently compared to hydrogen and methane in porous media, as also reported by Higgs2023. Despite the interfacial tension (IFT) of nitrogen (0.07 N m-1) and hydrogen (0.069 N m-1) being comparable at 10 MPa, the differences in viscosity significantly influence wetting-phase saturation at onset of drainage. This underscores the finding that at low fractional flows, the viscosity ratio plays a more pronounced role in gas displacement behavior. The fluid distribution in the fracture at stable fractional flow, = 0.6 as shown in Figure 7 (middle horizontal panel), indicates that nitrogen has the highest gas fraction (Sw = 0.76) in the fracture, followed by methane (Sw = 0.79), and then hydrogen (Sw = 0.83 ).
@@
At higher gas fractional flows ( = 1) the saturation converges for all gases as shown in Figure 6(b). At these higher gas fractional flows, the capillary number (Cat) decreases resulting in more connected gas which diminishes the impact of the viscosity ratio. Consequently, similar saturations are observed at = 1, representing endpoint saturation. Although the nitrogen case has a significantly higher Cat (1.2 x 10-6) than hydrogen and methane (7.0 x 10-6 and 1.5 x 10-6 respectively), the end point saturations remain consistent for all gases at Sw = 0.6. This suggests that the effect of capillary number on overall saturation is relatively minor under the tested conditions.
@@
The fluid distribution for the hydrogen-brine system in the fracture exhibits signs of intermittency at higher water saturations (lower ), as illustrated in Figure 8(a) for . Subtle differences in gas distribution are observed between two scans taken 30 minutes apart after reaching steady-state pressure. Several locations are identified by coloured circles: Circle 1 (yellow) marks a region where gas disappears in the subsequent scan, while Circle 2 (green) indicate location where gas reappears. These fluctuations in the gas phase correspond to capillary events and are also reflected in the measured differential pressure. In contrast, the nitrogen-brine system exhibits fewer fluctuations in saturation, suggesting a more stable flow pattern. A more detailed analysis of pressure fluctuations, intermittency, and their impact on fluid distribution will be discussed in Section 3.3. Along the length of the fracture, slice-averaged saturation reaches Sw = 1 for many slices, indicating a disconnected gas phase. Figure 8(b) presents the 1-D slice-averaged saturation profiles for these two subsequent scans revealing significant fluctuations in saturation within this period, emphasizing dynamic capillary processes occurring in the fracture. These processes are not fully captured in terms of intermediate gray values due to the scanner’s limited temporal resolution. As the gas fractional flow increases to = 1, the gas phase becomes more connected throughout the fracture, which reduces intermittency and enhances overall gas saturation as demonstrated by the 3D volume rendering of fluid distribution in Figure 7 (bottom horizontal panel). Intermittency significantly impacts relative permeability, resulting in notably low gas permeabilities, particularly for hydrogen and methane as discussed in Section 3.1.
@@
We note here that the observed intermittency may also be influenced by the dissolution of hydrogen and methane into the brine during the experiment, as the injected gas and brine were not pre-equilibrated, a phenomenon previously described by [Boon2022]. While bulk solubilities (hydrogen: 1.6 mg L-1, methane: 23 mg L-1, and nitrogen 19 mg L-1 [Kaye]) suggest hydrogen dissolution is minimal, but pore-scale dissolution kinetics governed by faster diffusion can still cause significant dissolution. However, the reappearance of gas in certain fracture locations indicates that intermittency is a genuine effect and not solely attributed to dissolution.
In addition to flow intermittency, geometric complexities may also contribute to the low gas permeabilities observed in our measurements. The dry scan (Figure 7 (left)) reveals secondary fractures branching from the central region of the main fracture, forming an intricate fracture network. As we discuss next, these geometric features influence fluid distribution and gas displacement efficiency, thereby affecting overall flow dynamics. The primary branch is highlighted in orange, while the secondary branch is shown in purple. The secondary branch is further subdivided into two regions based on aperture distribution, indicated by light blue and green circles. No gas invasion is observed near the inlet of the secondary branch (green region) for any gas or fractional flow, indicating a high capillary entry pressure due to the narrow aperture. At higher fractional flows ( = 0.6 and = 1) as shown in Figure 7 (middle and bottom horizontal panel, respectively), gas is present for all gas types in the secondary fracture near the outlet (ligh blue region), where the aperture is wider and better connected. However, at lower fractional flows (fg = 0.1) (Figure 7, top horizontal panel) particularly for low-viscosity gases like hydrogen and methane, gas does not invade this location, suggesting limited gas connectivity in fracture networks under these conditions.
@@ 3.3 Pressure analysis
The lower gas relative permeabilities for similar global water saturation suggest higher phase interference in the fracture. Saturation and fluid distribution in the fracture, based on scans taken 30 minutes apart, indicate the presence of intermittency, where Haines jumps and snap-off events occur. Since the scan duration is 16 minutes and these events appear to happen much more rapidly, we observed pressure data recorded at a much higher time resolution of 1 second. The fluctuations in differential pressure measurements from the transducer across the rock samples provide insights into the timescale and the fluctuations that correspond to snap-off and Haines jumps occurring during two-phase displacements [Spurin2022, Wang2023a]. Differential pressure measurements at over a 30-minute period for H2, CH4, and N2 are presented in Figure 9. These data reveal fluctuations over varied timescales for each gas, underscoring distinct dynamics associated with each gas phase. There are notable differences in the amplitude and frequency of these oscillations, particularly between H2 and CH4 on one hand and N2 on the other. The pressure amplitude of fluctuations is greater for hydrogen and methane compared to nitrogen, with the fluctuations occurring over timescales of approximately 150 seconds for hydrogen, 30 seconds for methane, and one second for nitrogen. The pressure fluctuations within the pores result in intermittent pore filling, occurring over a broad range of timescales. At different fractional flows, the timescales of these fluctuations vary, reflecting changes in the connectivity of the gas phase. The additional pressure data are provided in Supplementary Information S2.
High-speed imaging, such as synchrotron-based X-ray tomography, would offer valuable insights into the pore-scale dynamics of these processes. For example, Phillips2023 demonstrated that flow in rough fractures exhibits greater phase intermittency than in porous media, even during slow drainage. Additionally, Spurin2022 validated the correlation between pressure fluctuations and fluid distribution in porous media, highlighting the potential of pressure data as a tool for understanding intermittent flow behavior.
@@ 3.4 Sensitivity analysis
As discussed in previous sections, relative permeabilities at reservoir conditions in fractures are strongly influenced by intermittency, leading to significant fluctuations in pressure and saturation. To assess the sensitivity of the hydrogen-brine system under these conditions, we repeated the experiment under identical flow conditions. Overall, qualitative trends in the relative permeability-saturation curves remained consistent across both runs, as shown in Figure 10. The water permeabilities () were quantitatively similar in both runs. Gas permeability, while consistently low across runs, exhibited slight differences, likely due to minor pressure fluctuation variations. Since gas permeabilities are very low, even slight variations in pressure fluctuations can result in differences in the measured relative gas permeability. The pressure fluctuations for these two runs at are shown in Figure 10(b). The data clearly exhibits cyclic fluctuations, with the timescales of these fluctuations differing between the two runs: in the first run, the period was 0.5 minutes, while in the repeated run, the period extended to 2.5 minutes. These variations in fluctuation timescales may contribute to slight differences in the calculated relative gas permeabilities. Such differences could arise due to slight changes in fracture morphology, potentially occurring during pressurization, as the fracture is sensitive to applied stress. Over the course of running different experiments presented in this paper , the rock sample underwent pressurization-depressurization cycles, potentially impacting fracture morphology. This is reflected in the observed decrease in absolute permeability before the experimental runs, from 23 D to 8 D. Despite these slight variations, the overall behavior remains consistent across both experimental runs, confirming that our workflow is robust and validating the results presented in this study. The 1-D saturation profiles for these two runs also confirm it as shown in Supplementary Information S3.
While ideal statistical validation would involve multiple repetitions for each gas - brine pair, such replication was not feasible due to the significant safety and logistical challenges of working with hydrogen at 10 MPa. Nevertheless, the observed intermittency and capillary trapping behavior were reproducible across two hydrogen experiments and consistent with trends observed in methane, a gas with comparable viscosity and interfacial tension. In contrast, nitrogen, which is supercritical and an order of magnitude higher viscosity exhibited significantly higher relative permeability. These results are also consistent with prior literature as further discussed in Section 4.
@@ 4 Discussion
Direct measurements of relative permeability for hydrogen compared to other gases in fractured rock samples are currently absent from the literature. Here, we compare our findings with existing experimental data and numerical models for different fluid systems in rough fractures, as shown in Figure 11. Compared to previous studies, the hydrogen gas relative permeability in our system is notably lower, indicating interference effects that are stronger than predicted by existing models such as [Watanabe2015], Gong2021, and Bertels2001. The limited available experimental data were collected under different flow conditions and using different fluids, making direct comparisons challenging. The fluid pairs used in various studies from the literature are summarized in Supplementary Information S4. The nitrogen permeability aligns more closely with the model proposed by Gong2021; however, the impact of fracture geometry could not be fully resolved based on these comparisons. The nitrogen-brine system relative permeabilities were also compared to models from the literature, with detailed discussions provided in the Supplementary Information S4. In contrast, water permeability in our results aligns more closely with the model suggested by Piri2007. Importantly, fracture morphology also influences these differences, so while such comparisons are informative, they are not definitive. Controlled experiments and models that account for specific fracture morphology in hydrogen-brine systems are essential for more conclusive insights.
Notably, recent studies in porous media report endpoint relative permeabilities for hydrogen as low as 10-3 [Boon2022, Lysyy2022], which aligns with our fracture flow observations and highlights hydrogen’s reduced connectivity compared to methane and nitrogen. For instance, Boon2022 reported H2 - brine relative permeability in Berea sandstone to be an order of magnitude lower than CO2 - water and N2 - water systems in the same rock, as shown by Pini2013. This indicates stronger phase interference for low viscosity gases such as hydrogen. Our measured fracture relative permeabilities, while lower than values reported in other fracture studies, are therefore consistent with trends observed in hydrogen-specific porous media literature. Thus, we further compare our results with the hydrogen specific literature of porous media supporting the interpretation of the specificity of hydrogen flow compared to other gases investigated in the literature.
@@
The lower hydrogen gas permeability observed in our fracture study is consistent with trends seen in porous media studies in Berea sandstone [Lysyy2022, Boon2022] and Bentheimer sandstone [Higgs2023], shown in Figure 12. Despite marked differences in sample geometry and flow conditions, the results remain qualitatively comparable, as all experiments were conducted at low capillary numbers (Cat). In porous media, the relative permeability curves from Higgs2023 exhibit closer alignment for hydrogen, methane, and nitrogen. However, our fracture study reveals greater differences in both wetting- and non-wetting-phase permeabilities. This divergence can likely be attributed to the increased intermittency in the fracture, resulting from the higher pressure conditions (10 MPa) in our study, where the interfacial tension (IFT) between hydrogen and methane differs more significantly, and nitrogen exists in a supercritical state. In contrast, the experiments by Higgs2023, conducted at 20 bar, were performed under conditions where the fluid properties of these gases were more similar. Additionally, our study shows a much higher irreducible saturation, exemplifying the dependence of capillary-dominated relative permeability curves on flow rate (injection rate) and geometry (porous vs. fractured). When comparing our hydrogen permeability results to those from Boon2022, who also conducted high-pressure experiments (10 MPa) at 18∘C under similar conditions, hydrogen and water permeabilities were generally higher across most saturations for porous media, though the general trend remains consistent. These comparisons highlight the increased complexity of fluid interactions in fractures compared to those observed in porous media.
Low hydrogen gas relative permeability () directly limits its mobility, requiring higher injection pressures to meet target storage rates. Studies in porous media, including CO2 storage simulations [Oostrom2016, Boon2021] and recent hydrogen-specific modeling [Bo2023], show that low endpoint values lead to compact, flatter gas plumes with relatively uniform saturation and low gas saturations beyond the dry-out zone. While this reduces storage efficiency, it also spatially confines the gas plume, potentially lowering leakage risks and simplifying monitoring. Bo2023, in a porous reservoir context, further demonstrated that increasing injection or production rates does not necessarily enhance performance due to limited flow and pressure distribution caused by gas trapping. Using relative permeability data from other fluid systems can therefore introduce significant errors in predicting hydrogen behavior. Moreover, compensating for low mobility by injecting at higher rates could risk exceeding caprock integrity thresholds. These insights highlight the need to extend such simulation studies to fractured reservoirs, using laboratory-measured hydrogen relative permeability data under relevant flow conditions. Accurate characterization of in site-specific fracture systems is critical to avoid underestimating injectivity limitations or overestimating storage capacity, and to ensure a realistic balance between operational feasibility and storage efficiency.
@@ 5 Conclusions
While this study utilizes a single artificially fractured sample, which cannot capture the full spectrum of natural fracture heterogeneity (e.g., aperture distributions, mineral heterogeneity, or network connectivity), its primary objective was to compare fluid-specific behavior of hydrogen, methane, and nitrogen under as identical fracture conditions as possible. The observed differences in relative permeability, intermittency, and endpoint saturation are therefore attributable to intrinsic gas properties (e.g., viscosity, dissolution, interfacial tension) rather than geometric variability. Importantly, the sample features a realistic branching network (Figure 1(a)) and significant roughness, providing important insights into gas trapping and connectivity. Given the absence of high-pressure hydrogen relative permeability data for fractures, our work provides a foundational dataset to validate pore-scale hydrogen transport models, support discrete fracture network (DFN) simulations that upscale permeability based on aperture distributions [Hyman2015], and guide future experimental design across varied fracture morphologies.
In conclusion, this study provides critical insights into hydrogen behavior in fractured rock, supporting efforts to optimize underground hydrogen storage. By highlighting the limitations of using nitrogen as a proxy for hydrogen, our findings emphasize the need for hydrogen-specific research to develop effective storage strategies and safety protocols in subsurface reservoirs. In this study, we considered only drainage relative permeabilities. However, cyclic operations are likely to induce hysteresis, particularly in the gas phase, as observed by Boon2024 in porous sandstone. A comprehensive understanding of cyclic underground hydrogen storage (UHS) at field scale would therefore require future work to measure or simulate imbibition relative permeabilities. In addition to imbibition relative permeabilities, the future work should focus on measuring hydrogen-brine relative permeability in fracture samples with varied aperture sizes and roughness. Such studies will be valuable in understanding the impact of fracture geometry on flow behavior, ultimately enhancing the accuracy of upscaled fracture network models for reservoir-scale applications.